OpenStudy (anonymous):

Helppp, i know this is either combinations or permutations but i dont get it! :((( The local news station always sends a crew of 3 employees to cover important stories. The station has 3 reporters and 5 camera operators available. If a crew consists of 1 reporter and 2 camera operators, how many different such crews are possible?

OpenStudy (anonymous):

type "khan academy probability" into google. the best time investment you will ever make.

OpenStudy (anonymous):

i will, but do you know how to solve this?

OpenStudy (anonymous):

one sec, i'll give it a go.

OpenStudy (anonymous):

i know that its combinations, since order is not mentioned.

OpenStudy (anonymous):

actually it is the counting principle

OpenStudy (anonymous):

there are 3 possibilities for reporters and \(\frac{5\times 4}{2}=10\) possible ways to pick two of the five camera operators

OpenStudy (anonymous):

C(3,1) and C(5,2)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so the answer by the counting principle is \(3\times 10=30\)

OpenStudy (anonymous):

don't get married to the formulas they all come from the counting principle, so if you think in terms of "how many" you will get the answer

OpenStudy (anonymous):